### v1

@@ Abstract
Interval identification of parameters such as average treatment effects, average partial effects and welfare is particularly common when using observational data and experimental data with imperfect compliance due to the endogeneity of individuals’ treatment uptake. In this setting, a treatment or policy will typically become an object of interest to the researcher when it is either selected from the estimated set of best-performers or arises from a data-dependent selection rule. In this paper, we develop new inference tools for interval-identified parameters chosen via these forms of selection. We develop three types of confidence intervals for data-dependent and interval-identified parameters, discuss how they apply to several examples of interest and prove their uniform asymptotic validity under weak assumptions.
Keywords: Partial Identification, Post-Selection Inference, Conditional Inference, Uniform Validity, Treatment Choice.
@@ 1 Introduction
There is now a large and growing literature on partial identification of optimal treatments and policies under practically-relevant assumptions for observational data and experimental data with imperfect compliance (e.g., Stoye,, 2012; Kallus and Zhou,, 2021; Pu and Zhang,, 2021; D’Adamo,, 2021; Yata,, 2021; Han,, 2024 to name just a few). Interval identification of average treatment effects (ATEs), average partial effects and welfare is particularly common in these settings due to the endogeneity of individuals’ treatment uptake. In order to use the partial identification results for treatment or policy choice in practice, a researcher must typically estimate a set of best-performing treatments or policies from data. Consequently, a treatment or policy will typically become an object of interest to the researcher when it is either selected from the estimated set of best-performers or arises from a data-dependent selection rule. It is now well-known that selecting an object of interest from data invalidates standard inference tools (e.g., Andrews et al.,, 2024). The failure of standard inference tools after data-dependent selection is only compounded by the presence of partially-identified parameters.
@@
The current literature has not yet developed valid post-selection inference tools in partially-identified contexts, an important deficiency that the methods proposed in this paper aim to correct. The methods we propose here build upon the ideas of conditional and hybrid inference employed in various point-identified contexts by, e.g., Lee et al., (2016), Andrews et al., (2024) and McCloskey, (2024) to produce confidence intervals (CIs) for interval-identified parameters such as welfare or ATEs chosen from an estimated set or via a data-dependent selection rule. Although Andrews et al., (2023) also propose conditional and hybrid inference methods in the partial identification context of moment inequality models, they do not allow for data-dependent selection of objects of interest, one of the main focuses of the present paper. Finally, this paper directly relies upon results in the literature on interval identification of welfare, treatment effects and partial outcomes such as Manski, (1990), Balke and Pearl, (1997, 2011), Manski and Pepper, (2000), Han and Yang, (2024) and Han, (2024). We apply our inference methods to a general class of problems nesting these examples.
After sketching the ideas behind our inference methods in a simple example, we introduce the general inference framework to which our methods can be applied. We show that our general inference framework incorporates several problems of interest for data-dependent selection and treatment rules for parameters belonging to an identified set such as Manski bounds for average potential outcomes or ATEs, bounds on parameters derived from linear programming (e.g., Balke and Pearl,, 1997, 2011; Mogstad et al.,, 2018; Han,, 2024; Han and Yang,, 2024), bounds on welfare for treatment allocation rules that are partially identified by observational data (e.g., Stoye,, 2012; Kallus and Zhou,, 2021; Pu and Zhang,, 2021; D’Adamo,, 2021; Yata,, 2021) and bounds for dynamic treatment effects (e.g., Han,, 2024). We also show how to incorporate inference on parameters chosen via asymptotically optimal treatment choice rules (e.g., Christensen et al.,, 2023) into our general inference framework. Our framework can also be applied to settings where welfare is partially identified for other reasons (e.g., Cui and Han,, 2024).
@@
The second and third types of CIs we develop in this paper are designed for inference on parameters chosen by data-dependent selection rules for which the object of interest is uniquely determined by the selection rule. The projection CIs do not require knowledge of the selection rule to be asymptotically valid whereas the hybrid CI construction utilizes the particular form of a selection rule to improve upon the length properties of the projection CI. The conditional CIs are short for selections that occur with high probability but can become exceptionally long when this probability is small (see, e.g., Kivaranovic and Leeb,, 2021). Conversely, projection CIs are overly conservative when selection probabilities are high. Hybrid CIs interpolate the length properties of the conditional and projection CIs in order to attain good length properties regardless of the value selection probabilities take. In analogy with the conditional CIs, we formally show that both projection and hybrid CIs have are asymptotically valid in a uniform sense.
@@
We analyze the coverage and length properties of our proposed CIs in finite samples. Since, to our knowledge, these are the first uniformly valid CIs for data-dependent selections of partially-identified parameters, there are no existing CIs to which we can direclty compare. Nevertheless, since our CIs can also be used for inference on a priori chosen interval-identified parameters, we conduct a power comparison with one of the leading methods for inference on a partially-identified parameter. In particular, we compare the power of a test implied by our hybrid CIs to the power of the hybrid test of Andrews et al., (2023), a test that applies to a general class of moment-inequality models that is also based upon a (different) hybrid between conditional and projection-based inference. Encouragingly, the power of the test implied by our hybrid CI is quite competitive even in this environment for which it was not designed. We also find that the finite-sample coverage of all of our CIs is approximately correct in a simple Manski bound example. Finally, we analyze the length tradeoffs between the three different CIs across different DGPs, finding the hybrid CI to perform best overall.
@@
The remainder of this paper is structured as follows. Section 2 sketches the ideas behind our general CI constructions in the context of a simple Manski bound example. Section 3 lays down the general high-level inference framework we are interested in while Section 4 details how the general framework applies in several different examples. Sections 5 and 6 then detail the various CI constructions in the general setting. The final two sections of the paper, Sections 7 and 8, are devoted to finite-sample comparisons of the properties of the different CIs in the context of a simple Manski bound example. Mathematical proofs are relegated to a Technical Appendix at the end of the paper.
@@ 2 Basic Ideas: Inference with Manski Bounds Example
Consider a binary outcome of interest $Y$, a binary treatment indicator $D$ and a binary treatment assignment $Z$. Furthermore, let $Y(0)$ and $Y(1)$ denote potential outcomes under no treatment ($D=0$) and treatment ($D=1$). Assuming $E[Y(d)|Z]=E[Y(d)]$, Manski, (1990) shows that we can bound the average potential outcomes $W(d)=\mathbb{E}[Y(d)]$ in the absence and presence of treatment as follows:
@@
for $d=0,1$ and $p^{ydz}=Pr(Y=y,D=d|Z=z)$. Given (2.1) it is natural to estimate the set of best-performing options $\mathcal{D}^{*}$, as a subset of the two options of treatment and no treatment, from an observed dataset of outcomes, treatments and treatment assignments $\{(Y_{i},D_{i},Z_{i})\}_{i=1}^{n}$ as those that are estimated to be undominated:
@@
where $\widehat{L}(d)\equiv\max\{\hat{p}^{1d0},\hat{p}^{1d1}\}$ and $\widehat{U}(d)\equiv\min\{1-\hat{p}^{0d0},1-\hat{p}^{0d1}\}$ with $\hat{p}^{ydz}$ being an empirical estimate of $p^{ydz}=Pr(Y=y,D=d|Z=z)$.
@@ 2.1 Why Does Standard Inference Fail?
Reason 1. is easy to see since $\widehat{L}(d)$ is the maximum of two normally distributed random variables in large samples when $d$ is chosen a priori. To better understand reason 2., note that the distribution of $\widehat{L}(d)$ given $d\in\widehat{\mathcal{D}}$ is the conditional distribution of the maximum of two normally distributed random variables given that the minimum of two other normally distributed random variables, $\widehat{U}(d)\equiv\min\{1-\hat{p}^{0d0},1-\hat{p}^{0d1}\}$, exceeds the maximum of yet another set of two normally distributed random variables, $\widehat{L}(d^{\prime})=\max\{\hat{p}^{1d^{\prime}0},\hat{p}^{1d^{\prime}1}\}$ for $d^{\prime}\neq d$. Unconditionally, $\widehat{L}(\hat{d})$ for any data-dependent choice of $\hat{d}\in\widehat{\mathcal{D}}$ is distributed as a mixture of the distributions of $\widehat{L}(0)$ and $\widehat{L}(1)$, neither of which are themselves normally distributed.
@@ 2.2 Conditional Confidence Intervals
Suppose that a researcher’s interest in inference on $L(d)$ only arises when $d$ is estimated to be in the set of best-performing options, viz., $d\in\widehat{\mathcal{D}}$. In such a case, we are interested in a probabilisitic lower bound for $L(d)$ that is approximately valid across repeated samples for which $d\in\widehat{\mathcal{D}}$, i.e., we would like to form a conditionally valid lower bound $\widehat{L}(d)_{\alpha}^{C}$ such that [^4] See Andrews et al., (2024) for an extensive discussion of when conditional vs unconditional validity is desirable for inference after selection.
@@
for some $\alpha\in(0,1)$ in large samples. To do so we characterize the conditional distribution of $\widehat{L}(d)$. Specifically, let $\hat{j}_{L}(d)=\operatorname*{argmax\;}_{j\in\{0,1\}}\hat{p}^{1dj}$. Lemma 5.1 of Lee et al., (2016) implies
$\widehat{L}(d)\left|\left\{d\in\widehat{\mathcal{D}},\hat{j}_{L}(d)=j_{L}^{*}\right\}\right.\sim\xi_{L}\left|\left\{\widehat{\mathcal{L}}_{L}(Z_{L})\leq\xi_{L}\leq\widehat{\mathcal{U}}_{L}(Z_{L})\right\}\right.,$ (2.3)
for some known functions ^$\widehat{\mathcal{L}}_{L}(\cdot)$ and $\widehat{\mathcal{U}}_{L}(\cdot)$, where $\xi_{L}\sim\mathcal{N}(p^{1dj_{L}^{*}},\operatorname*{Var}(\hat{p}^{1dj_{L}^{*}}))$ in large samples and $Z_{L}$ is a sufficient statistic for the nuisance parameter $p=(p^{100},p^{010},p^{110},p^{101},p^{011},p^{111})^{\prime}$ that is asymptotically independent of $\xi_{L}$. Using results in Pfanzagl, (1994), the characterization in (2.3) permits the straightforward computation of a conditionally quantile-unbiased estimator for $p^{1d\hat{j}_{L}(d)}$, $\widehat{L}(d)_{\alpha}^{C}$, satisfying
@@
Finally, an analogous construction to that described above for producing a probabilistic lower bound for $L(d)$ produces a conditionally valid probabilistic upper bound $\widehat{U}(d)_{1-\alpha}^{C}$ for $U(d)$ that satisfies
@@ 2.3 Unconditional Confidence Intervals
First, we can form a probabilistic lower bound for $L(\hat{d})$ by projecting a one-sided rectangular simultaneous confidence lower bound for all possible values $L(\hat{d})$ can take: $\widehat{L}(\hat{d})_{\alpha}^{P}=\widehat{L}(\hat{d})-\sqrt{\widehat{\Sigma}_{L,2d+1+\hat{j}_{L}(d)}}\hat{c}_{1-\alpha,L},$ where $\hat{c}_{1-\alpha,L}$ is the $1-\alpha$ quantile of $\max_{i}\hat{\zeta}_{i}/\sqrt{\widehat{\Sigma}_{L,i}}$ for $\hat{\zeta}\sim\mathcal{N}(0,\widehat{\Sigma}_{L})$, $\widehat{\Sigma}_{L}$ is a consistent estimator of $\Sigma_{L}=\operatorname*{Var}(\hat{p}^{100},\hat{p}^{101},\hat{p}^{110},\hat{p}^{111})$ and $\mathcal{M}_{i}$ denotes the $i^{th}$ element of the main diagonal of the matrix $\mathcal{M}$. Since
$\mathbb{P}\left((p^{100},p^{101},p^{110},p^{111})\geq(\hat{p}^{100},\hat{p}^{101},\hat{p}^{110},\hat{p}^{111})-\sqrt{\operatorname{Diag}(\widehat{\Sigma}_{L})}\hat{c}_{1-\alpha,L}\right)=1-\alpha,$
in large samples, $\mathbb{P}(p^{1\hat{d}\hat{j}_{L}(\hat{d})}\geq\widehat{L}(\hat{d})_{\alpha}^{P})\geq 1-\alpha$ and (2.7) holds for $\widehat{L}(\hat{d})_{\alpha}^{U}=\widehat{L}(\hat{d})_{\alpha}^{P}$ becasue $L(\hat{d})\geq p^{1\hat{d}\hat{j}_{L}(\hat{d})}$. However, $\widehat{L}(\hat{d})_{\alpha}^{P}$ suffers from a converse drawback to that of $\widehat{L}(\hat{d})_{\alpha}^{C}$: it is unnecessarily conservative when $\hat{d}=d$ is chosen with high probability (see e.g., Andrews et al.,, 2024 and McCloskey,, 2024).
We propose a second probabilisitc lower bound for $L(\hat{d})$ that combines the complementary strengths of $\widehat{L}(d)_{\alpha}^{C}$ and $\widehat{L}(\hat{d})_{\alpha}^{P}$. Construction of this hybrid lower bound $\widehat{L}(\hat{d})_{\alpha}^{H}$ proceeds analogously to the construction of $\widehat{L}(d)_{\alpha}^{C}$ after adding the additional condition $\{p^{1\hat{d}\hat{j}_{L}(\hat{d})}\geq\widehat{L}(\hat{d})_{\beta}^{P}\}$ for $\beta<\alpha$ to the conditioning event and instead computing a conditionally quantile-unbiased estimator for $p^{1\hat{d}\hat{j}_{L}(\hat{d})}$, $\widehat{L}(\hat{d})_{\alpha}^{H}$, satisfying
@@
in large samples, where $d^{*}$ is any realized value of the random variable $\hat{d}$. Imposing this additional condition in the formation of the hybrid bound ensures that $\widehat{L}(\hat{d})_{\alpha}^{H}$ is always greater than $\widehat{L}(\hat{d})_{\beta}^{P}$, limiting its worst-case performance relative to $L(\hat{d})_{\beta}^{P}$ when $\mathbb{P}(\hat{d}=d^{*})$ is small. On the other hand, when $\mathbb{P}(\hat{d}=d^{*})$ is large, the additional condition $\{p^{1\hat{d}\hat{j}_{L}(\hat{d})}\geq\widehat{L}(\hat{d})_{\beta}^{P}\}$ is far from binding with high probability so that $\widehat{L}(\hat{d})_{\alpha}^{H}$ becomes very close to $\widehat{L}(d)_{(\alpha-\beta)/(1-\beta)}^{C}$. In this case, $\widehat{L}(d)_{(\alpha-\beta)/(1-\beta)}^{C}$ is close to the naive lower bound based upon the normal distribution $\widehat{L}(\hat{d})-\sqrt{\operatorname*{Var}(\hat{p}^{1dj_{L}^{*}})}z_{(1-\alpha)/(1-\beta)}$ because the truncation bounds in (2.3) are very wide (Proposition 3 in Andrews et al.,, 2024).
@@
$\mathbb{P}\left(\left.L(\hat{d})\geq\widehat{L}(\hat{d})_{\alpha}^{H}\right|\hat{d}=d^{*},p^{1\hat{d}\hat{j}_{L}(\hat{d})}\geq\widehat{L}(\hat{d})_{\beta}^{P}\right)$
$\geq\mathbb{P}\left(\left.p^{1\hat{d}\hat{j}_{L}(\hat{d})}\geq\widehat{L}(\hat{d})_{\alpha}^{H}\right|\hat{d}=d^{*},p^{1\hat{d}\hat{j}_{L}(\hat{d})}\geq\widehat{L}(\hat{d})_{\beta}^{P}\right)=\frac{1-\alpha}{1-\beta}$
@@
By similar reasoning to that used for the conditional confidence intervals in Section 2.2 above, $\widehat{L}(\hat{d})_{\alpha}^{H}$ is not very conservative as a probabilistic lower bound for $L(\hat{d})$. The researcher’s choice of $\beta\in(0,\alpha)$ trades off the performance of $\widehat{L}(\hat{d})_{\alpha}^{H}$ across scenarios for which $\mathbb{P}(\hat{d}=d^{*})$ is large and small with a small $\beta$ corresponding to better performance when $\mathbb{P}(\hat{d}=d^{*})$ is large. See McCloskey, (2024) for an in-depth discussion of these tradeoffs. We recommend $\beta=\alpha/10$.

### v2

@@ Abstract
Interval identification of parameters such as average treatment effects, average partial effects and welfare is particularly common when using observational data and experimental data with imperfect compliance due to the endogeneity of individuals’ treatment uptake. In this setting, the researcher is typically interested in a treatment or policy that is either selected from the estimated set of best-performers or arises from a data-dependent selection rule. In this paper, we develop new inference tools for interval-identified parameters chosen via these forms of selection. We develop three types of confidence intervals for data-dependent and interval-identified parameters, discuss how they apply to several examples of interest and prove their uniform asymptotic validity under weak assumptions.
Keywords: Partial Identification, Post-Selection Inference, Selective Inference, Conditional Inference, Uniform Validity, Treatment Choice.
@@ 1 Introduction
There is now a large and growing literature on partial identification of optimal treatments and policies under practically-relevant assumptions for observational data and experimental data with imperfect compliance (e.g., Stoye,, 2012; Kallus and Zhou,, 2021; Pu and Zhang,, 2021; D’Adamo,, 2021; Yata,, 2021; Han,, 2024 to name just a few). Interval identification of average treatment effects (ATEs), average partial effects and welfare is particularly common in these settings due to the endogeneity of individuals’ treatment uptake. In order to use the partial identification results for treatment or policy choice in practice, a researcher must typically estimate a set of best-performing treatments or policies from data. Consequently, the researcher is typically interested in a treatment or policy that is either selected from the estimated set of best-performers or arises from a data-dependent selection rule. It is now well-known that selecting an object of interest from data invalidates standard inference tools (e.g., Andrews et al.,, 2024). The failure of standard inference tools after data-dependent selection is only compounded by the presence of partially-identified parameters.
@@
The current literature has not yet developed valid post-selection inference tools in partially-identified contexts, an important deficiency that the methods proposed in this paper aim to correct. The methods we propose here build upon the ideas of conditional and hybrid inference employed in various point-identified contexts by, e.g., Lee et al., (2016), Fithian et al., (2017), Tibshirani et al., (2018), Andrews et al., (2024) and McCloskey, (2024) to produce confidence intervals (CIs) for interval-identified parameters such as welfare or ATEs chosen from an estimated set or via a data-dependent selection rule. Although Andrews et al., (2023) also propose conditional and hybrid inference methods in the partial identification context of moment inequality models, they do not allow for data-dependent selection of objects of interest, one of the main focuses of the present paper. Finally, this paper directly relies upon results in the literature on interval identification of welfare, treatment effects and partial outcomes such as Manski, (1990), Balke and Pearl, (1997, 2011), Manski and Pepper, (2000), Mogstad et al., (2018), Han and Yang, (2024) and Han, (2024). We apply our inference methods to a general class of problems nesting these examples.
After sketching the ideas behind our inference methods in a simple example, we introduce the general inference framework to which our methods can be applied. We show that our general inference framework incorporates several problems of interest for data-dependent selection and treatment rules for parameters belonging to an identified set such as Manski bounds for average potential outcomes or ATEs, bounds on parameters derived from linear programming (e.g., Balke and Pearl,, 1997, 2011; Mogstad et al.,, 2018; Han,, 2024; Han and Yang,, 2024), bounds on welfare for treatment allocation rules that are partially identified by observational data (e.g., Stoye,, 2012; Kallus and Zhou,, 2021; Pu and Zhang,, 2021; D’Adamo,, 2021; Yata,, 2021) and bounds for dynamic treatment effects (e.g., Han,, 2024). We also show how to incorporate inference on parameters chosen via asymptotically optimal treatment choice rules (e.g., Christensen et al.,, 2023) into our general inference framework. Our framework can also be applied to settings where welfare is partially identified for reasons other than treatment endogeneity (e.g., Ishihara and Kitagawa,, 2021; Adjaho and Christensen,, 2022; Ben-Michael et al.,, 2021; Cui and Han,, 2024).
@@
The second and third types of CIs we develop in this paper are designed for inference on parameters chosen by data-dependent selection rules for which the object of interest is uniquely determined by the selection rule. The projection CIs do not require knowledge of the selection rule to be asymptotically valid, whereas the hybrid CI construction utilizes the particular form of a selection rule to improve upon the length properties of the projection CI. The conditional CIs are short for selections that occur with high probability but can become exceptionally long when this probability is small (see, e.g., Kivaranovic and Leeb,, 2021). Conversely, projection CIs are overly conservative when selection probabilities are high. Although conditional CIs can be used in this setting, hybrid CIs interpolate the length properties of the conditional and projection CIs in order to attain good length properties regardless of the value selection probabilities take. In analogy with the conditional CIs, we formally show that both projection and hybrid CIs are asymptotically valid in a uniform sense.
@@
We analyze the coverage and length properties of our proposed CIs in finite samples. Since, to our knowledge, these are the first uniformly valid CIs for data-dependent selections of partially-identified parameters, there are no existing CIs to which we can directly compare. Nevertheless, since our CIs can also be used for inference on a priori chosen interval-identified parameters, we conduct a power comparison with one of the leading methods for inference on a partially-identified parameter. In particular, we compare the power of the test implied by our hybrid CIs to the power of the hybrid test of Andrews et al., (2023), a test that applies to a general class of moment-inequality models that is also based upon a (different) hybrid between conditional and projection-based inference. Encouragingly, the power of the test implied by our hybrid CI is quite competitive even in this environment for which it was not designed. We also find that the finite-sample coverage of all of our CIs is approximately correct in a simple Manski bound example. Finally, we analyze the length tradeoffs between the three different CIs across different DGPs, finding the hybrid CI to perform best overall.
@@
The remainder of this paper is structured as follows. Section 2 sketches the ideas behind our general CI constructions in the context of a simple Manski bound example. Section 3 lays down the general high-level inference framework we are interested in, while Section 4 details how the general framework applies in several different examples. Section 5 then details the various CI constructions in the general setting. Sections 6 and 7 are devoted to finite-sample comparisons of the properties of the different CIs in the context of a simple Manski bound example. The final section, Section 8, contains an empirical application where we apply our procedures to dynamic policies of schooling and post-school training. Appendix A contains additional examples that fit our general framework that are not covered in Section 4, while Appendix B contains additional simulation results corresponding to the dynamic treatment regime example detailed in Section 8. Mathematical proofs are relegated to a Technical Appendix at the end of the paper.
@@ 2 Basic Ideas: Inference with Manski Bounds Example
We first provide a simple example to illustrate our proposed methods. Consider a binary outcome of interest $Y$, a binary treatment indicator $D$ and a binary treatment assignment $Z$. Furthermore, let $Y(1)$ and $Y(0)$ denote potential outcomes under treatment ($D=1$) and no treatment ($D=0$). Assuming $\mathbb{E}[Y(d)|Z]=\mathbb{E}[Y(d)]$, Manski, (1990) shows that we can bound the average potential outcomes $W(d)=\mathbb{E}[Y(d)]$ in the absence and presence of treatment as follows:
@@
for $d=0,1$ and $p^{ydz}\equiv Pr(Y=y,D=d|Z=z)$. Given (2.1) it is natural to define the set of best-performing options $\mathcal{D}^{*}$, as a subset of the two options of treatment and no treatment, to be those that are undominated options. From an observed dataset of outcomes, treatments and treatment assignments $\{(Y_{i},D_{i},Z_{i})\}_{i=1}^{n}$, such a set can be estimated as:
@@
where $\widehat{L}(d)\equiv\max\{\hat{p}^{1d0},\hat{p}^{1d1}\}$ and $\widehat{U}(d)\equiv\min\{1-\hat{p}^{0d0},1-\hat{p}^{0d1}\}$ with $\hat{p}^{ydz}$ being an empirical estimate of the fitted probability $p^{ydz}\equiv\mathbb{P}(Y=y,D=d|Z=z)$.
@@ 2.1 Why Does Standard Inference Fail?
Reason 1. is easy to see since $\widehat{L}(d)$ is the maximum of two normally distributed random variables in large samples when $d$ is chosen a priori. To better understand reason 2., note that the distribution of $\widehat{L}(d)$ given $d\in\widehat{\mathcal{D}}$ is the conditional distribution of the maximum of two normally distributed random variables given that the minimum of two other normally distributed random variables, $\widehat{U}(d)\equiv\min\{1-\hat{p}^{0d0},1-\hat{p}^{0d1}\}$, exceeds the maximum of yet another set of two normally distributed random variables, $\widehat{L}(d^{\prime})=\max\{\hat{p}^{1d^{\prime}0},\hat{p}^{1d^{\prime}1}\}$ for $d^{\prime}\neq d$. Unconditionally, $\widehat{L}(\hat{d})$ for any data-dependent choice of $\hat{d}$ is distributed as a mixture of the distributions of $\widehat{L}(0)$ and $\widehat{L}(1)$, neither of which are themselves normally distributed.
@@ 2.2 Conditional Confidence Intervals
Suppose that a researcher’s interest in inference on $L(d)$ only arises when $d$ is estimated to be in the set of best-performing options, viz., $d\in\widehat{\mathcal{D}}$. In such a case, we are interested in a probabilistic lower bound for $L(d)$ that is approximately valid across repeated samples for which $d\in\widehat{\mathcal{D}}$, i.e., we would like to form a conditionally valid lower bound $\widehat{L}(d)_{\alpha}^{C}$ such that [^4] See Andrews et al., (2024) for an extensive discussion of when conditional vs unconditional validity is desirable for inference after selection.
@@
for some $\alpha\in(0,1)$ in large samples. To do so we characterize the conditional distribution of $\widehat{L}(d)$. Specifically, let $\hat{j}_{L}(d)\equiv\operatorname*{argmax\;}_{j\in\{0,1\}}\hat{p}^{1dj}$ be the value of $Z$ at which the maximum between the two estimated probabilities is achieved. Then $\widehat{L}(d)=\hat{p}^{1d\hat{j}_{L}(d)}$. Also, since the conditioning event $\{d\in\widehat{\mathcal{D}},\hat{j}_{L}(d)=j_{L}^{*}\}$ can be written as a polyhedron in $\hat{p}$ and $\widehat{L}(d)$ is equal to an element of $\hat{p}$, Lemma 5.1 of Lee et al., (2016) implies
$\widehat{L}(d)\left|\left\{d\in\widehat{\mathcal{D}},\hat{j}_{L}(d)=j_{L}^{*}\right\}\right.\sim\hat{p}^{1dj_{L}^{*}}\left|\left\{\widehat{\mathcal{L}}_{L}(\mathcal{Z}_{L})\leq\hat{p}^{1dj_{L}^{*}}\leq\widehat{\mathcal{U}}_{L}(\mathcal{Z}_{L})\right\}\right.$ (2.3)
for some known functions ^$\widehat{\mathcal{L}}_{L}(\cdot)$ and $\widehat{\mathcal{U}}_{L}(\cdot)$, where $\hat{p}^{1dj_{L}^{*}}\sim\mathcal{N}(p^{1dj_{L}^{*}},\operatorname*{Var}(\hat{p}^{1dj_{L}^{*}}))$ in large samples and $\mathcal{Z}_{L}$ is a sufficient statistic for the nuisance parameter $p=(p^{100},p^{010},p^{110},p^{101},p^{011},p^{111})^{\prime}$ that is asymptotically independent of $\hat{p}^{1dj_{L}^{*}}$. Using results in Pfanzagl, (1994), the characterization in (2.3) permits the straightforward computation of a conditionally quantile-unbiased estimator for $p^{1d\hat{j}_{L}(d)}=p^{1dj_{L}^{*}}$, since the latter is equal to the mean of the underlying normally distributed random variable $\hat{p}^{1dj_{L}^{*}}$ that is subject to truncation. Denoting this quantile-unbiased estimator as $\widehat{L}(d)_{\alpha}^{C}$, we have
@@
Finally, a construction analogous to that described above for producing a probabilistic lower bound for $L(d)$ produces a conditionally valid probabilistic upper bound $\widehat{U}(d)_{1-\alpha}^{C}$ for $U(d)$ that satisfies
@@ 2.3 Unconditional Confidence Intervals
First, we can form a probabilistic lower bound for $L(\hat{d})$ by projecting a one-sided rectangular simultaneous confidence lower bound for all possible values $L(\hat{d})$ can take: $\widehat{L}(\hat{d})_{\alpha}^{P}\equiv\widehat{L}(\hat{d})-\hat{c}_{1-\alpha,L}\sqrt{\widehat{\Sigma}_{L,2\hat{d}+1+\hat{j}_{L}(\hat{d})}},$ where $\hat{c}_{1-\alpha,L}$ is the $1-\alpha$ quantile of $\max_{i}\hat{\zeta}_{i}/\sqrt{\widehat{\Sigma}_{L,i}}$ for $\hat{\zeta}\sim\mathcal{N}(0,\widehat{\Sigma}_{L})$, $\widehat{\Sigma}_{L}$ is a consistent estimator of $\Sigma_{L}\equiv\operatorname*{Var}(\hat{p}^{100},\hat{p}^{101},\hat{p}^{110},\hat{p}^{111})$ and $\Upsilon_{i}$ denotes the $i^{th}$ element of the main diagonal of any square matrix $\Upsilon$. Here, the maximum is taken to guarantee simultaneous coverage of all possible values of $L(\hat{d})$. Since $p^{1\hat{d}\hat{j}_{L}(\hat{d})}\in\{\hat{p}^{100},\hat{p}^{101},\hat{p}^{110},\hat{p}^{111}\}$ with probability one,
$\mathbb{P}\left(p^{1\hat{d}\hat{j}_{L}(\hat{d})}\geq\widehat{L}(\hat{d})_{\alpha}^{P}\right)=\mathbb{P}\left(p^{1\hat{d}\hat{j}_{L}(\hat{d})}\geq\hat{p}^{1\hat{d}\hat{j}_{L}(\hat{d})}-\hat{c}_{1-\alpha,L}\sqrt{\widehat{\Sigma}_{L,2\hat{d}+1+\hat{j}_{L}(\hat{d})}}\right)$
$\geq\mathbb{P}\left((p^{100},p^{101},p^{110},p^{111})\geq(\hat{p}^{100},\hat{p}^{101},\hat{p}^{110},\hat{p}^{111})-\hat{c}_{1-\alpha,L}\sqrt{\operatorname{Diag}(\widehat{\Sigma}_{L})}\right)=1-\alpha$
in large samples and (2.7) holds for $\widehat{L}(\hat{d})_{\alpha}^{U}=\widehat{L}(\hat{d})_{\alpha}^{P}$ because $L(\hat{d})\geq p^{1\hat{d}\hat{j}_{L}(\hat{d})}$. However, $\widehat{L}(\hat{d})_{\alpha}^{P}$ suffers from a converse drawback to that of $\widehat{L}(\hat{d})_{\alpha}^{C}$: it is unnecessarily conservative when $\hat{d}=d$ is chosen with high probability (see e.g., Andrews et al.,, 2024 and McCloskey,, 2024).
We propose a second probabilistic lower bound for $L(\hat{d})$ that combines the complementary strengths of $\widehat{L}(d)_{\alpha}^{C}$ and $\widehat{L}(\hat{d})_{\alpha}^{P}$. Construction of this hybrid lower bound $\widehat{L}(\hat{d})_{\alpha}^{H}$ proceeds analogously to the construction of $\widehat{L}(d)_{\alpha}^{C}$ after adding the additional condition $\{p^{1\hat{d}\hat{j}_{L}(\hat{d})}\geq\widehat{L}(\hat{d})_{\beta}^{P}\}$ for $\beta<\alpha$ to the conditioning event and instead computing a conditionally quantile-unbiased estimator for $p^{1\hat{d}\hat{j}_{L}(\hat{d})}$, denoted as $\widehat{L}(\hat{d})_{\alpha}^{H}$, satisfying
@@
in large samples, where $d^{*}$ is any realized value of the random variable $\hat{d}$. Imposing this additional condition in the formation of the hybrid bound ensures that $\widehat{L}(\hat{d})_{\alpha}^{H}$ is always greater than $\widehat{L}(\hat{d})_{\beta}^{P}$, limiting its worst-case performance relative to $L(\hat{d})_{\beta}^{P}$ when $\mathbb{P}(\hat{d}=d^{*})$ is small. On the other hand, when $\mathbb{P}(\hat{d}=d^{*})$ is large, the additional condition $\{p^{1\hat{d}\hat{j}_{L}(\hat{d})}\geq\widehat{L}(\hat{d})_{\beta}^{P}\}$ is far from binding with high probability so that $\widehat{L}(\hat{d})_{\alpha}^{H}$ becomes very close to $\widehat{L}(d)_{(\alpha-\beta)/(1-\beta)}^{C}$. In this case, $\widehat{L}(d)_{(\alpha-\beta)/(1-\beta)}^{C}$ is close to the naive lower bound based upon the normal distribution $\widehat{L}(\hat{d})-z_{(1-\alpha)/(1-\beta)}\sqrt{\operatorname*{Var}(\hat{p}^{1dj_{L}^{*}})}$ because the truncation bounds in (2.3) are very wide (Proposition 3 in Andrews et al.,, 2024).
@@
$\mathbb{P}\left(\left.L(\hat{d})\geq\widehat{L}(\hat{d})_{\alpha}^{H}\right|\hat{d}=d^{*},p^{1\hat{d}\hat{j}_{L}(\hat{d})}\geq\widehat{L}(\hat{d})_{\beta}^{P}\right)\geq\mathbb{P}\left(\left.p^{1\hat{d}\hat{j}_{L}(\hat{d})}\geq\widehat{L}(\hat{d})_{\alpha}^{H}\right|\hat{d}=d^{*},p^{1\hat{d}\hat{j}_{L}(\hat{d})}\geq\widehat{L}(\hat{d})_{\beta}^{P}\right)=\frac{1-\alpha}{1-\beta}$
@@
By similar reasoning to that used for the conditional CIs in Section 2.2 above, $\widehat{L}(\hat{d})_{\alpha}^{H}$ is not very conservative as a probabilistic lower bound for $L(\hat{d})$. The researcher’s choice of $\beta\in(0,\alpha)$ trades off the performance of $\widehat{L}(\hat{d})_{\alpha}^{H}$ across scenarios for which $\mathbb{P}(\hat{d}=d^{*})$ is large and small with a small $\beta$ corresponding to better performance when $\mathbb{P}(\hat{d}=d^{*})$ is large. See McCloskey, (2024) for an in-depth discussion of these tradeoffs. We recommend $\beta=\alpha/10$.